OpenStudy (anonymous):

write an equation in standard form using only integers for the line described. The line through (5,4),parallel to y=-2/7x +1

OpenStudy (anonymous):

If you need to be parallel to that other line, then you'll use the SAME slope as that other line. The slope of that other line is -2/7 because that's the coefficient on the x-term. Then you use the formula \[y-y _{1}=m(x-x _{1)}\] plugging in your values. \[y-4=-2/7(x-5)\]

OpenStudy (anonymous):

Lastly, you'll have to clean that up a bit - distribute the -2/7 to eliminate the parentheses. Then because the problem said only to use integers in your answer you have to remove the fractions by multiplying EVERY term by 7.

OpenStudy (anonymous):

Oh and it also says to leave the answer in standard form....so you need to get the x and the y-terms on one side. \[y-4=-2/7x+5/7\] \[y \times7 - 4\times7=-2/7x \times7+5/7\times7\] \[7y-28=-2x+5\] etc.